OpenStudy (destinykiara99):

Determine the domain of the function. f as a function of x is equal to four divided by x squared.

OpenStudy (destinykiara99):

Answer Choices: x ≥ 0 All real numbers except 0 All real numbers except 3 All real numbers

OpenStudy (xapproachesinfinity):

So the function is \[f(x)=\frac{4}{\sqrt{x}}\]

OpenStudy (destinykiara99):

\[f(x)=4/x^2\]

OpenStudy (xapproachesinfinity):

oh yeah miss reading

OpenStudy (xapproachesinfinity):

ok so all you need to remember is that a domain of a function is where it is defined

OpenStudy (xapproachesinfinity):

what I'm saying is values for which the function does not exit have to be excluded

OpenStudy (xapproachesinfinity):

so in your case what values makes f(x) undefined?

OpenStudy (destinykiara99):

I honestly have no idea, I don't understand any of this.

OpenStudy (xapproachesinfinity):

well ok! let me ask you this can we divide by 0? is it legitimate ?

OpenStudy (xapproachesinfinity):

ok take for instance can we do 4/0 ?

OpenStudy (destinykiara99):

No

OpenStudy (xapproachesinfinity):

then we can say that for 0 the function is undefined

OpenStudy (xapproachesinfinity):

So we say the domain is all real numbers except 0

OpenStudy (xapproachesinfinity):

because all other values are allowed but 0 no

OpenStudy (xapproachesinfinity):

do you get it now?

OpenStudy (destinykiara99):

Yes I do, thank you!

OpenStudy (xapproachesinfinity):

no problem anytime you have a fraction you need to set denominator = 0 to find the values that make it 0 that's how you get the domain for rational functions